# Equitable Multi-Task Learning for AI-RANs

### v1

@@ I Introduction
Still, MTL on its own has already shown significant promise across various applications. For instance, it was found to improve RAN-related tasks such as carrier prediction, resource management, modulation and signal classification, as well as radio environment and mobility characterization [15, 1, 10, 28, 7]. Further, MTL has been used for multi-objective network optimization, collaborative control and multi-task edge inference, demonstrating that shared model representations can enhance learning [20, 24]. Clearly, a plausible next step is to use MTL for dynamic tasks of users.
@@
The operation of this system is modeled as an OWO fairness problem that we tackle with a two-layer learning algorithm based on the theory of Online Convex Optimization (OCO). At the higher layer (outer-loop), the system learns to initialize the model at the start of each round; where a good initialization helps the model to adapt fast to the forthcoming jobs. At the lower layer (inner-loop) the system updates the model with the feedback [^1] The feedback can be created using a small testing dataset as in [4, 8], or might be directly observable in case of measurable tasks [7], or even obtained through standard self/un-supervised learning metrics. the users provide after each slot’s jobs are executed. The algorithm offers optimality guarantees or, formally, zero fairness regret: it achieves the same performance as a model that would have been selected by a clairvoyant oracle, and this holds for any possible sequence of jobs and tasks. At the same time, the algorithm has minimal computing and storage requirements, which is in stark contrast to other MTL methods that calculate and store as many gradients as the tasks. Our extensive performance evaluation using both (convex) kernel regression and (non-convex) deep learning experiments, verifies the practical gains of the algorithm over several benchmarks and state-of-the-art MTL solutions.
@@ II System Model and Problem Statement
Model. We consider an AI-RAN system serving a set of users $\mathcal{K}$, who create ML tasks in a streaming fashion. The system operates in mixed-time scale with $m$ slots per round $t\!\leq\!T$. During each round $t$, every user $k\in\mathcal{K}$ needs to perform a task, denoted $(k,t)$, that involves $m$ sequential jobs (one per slot). The system supports the users through a shared model which is trained dynamically with data from all tasks. For instance, consider a mobile video analytics service running at the edge [9]. Each user $k$ corresponds to a mobile device with the task of object detection through a video stream. A single task $(k,t)$ then represents one usage session of the service under a particular context, e.g., lighting and weather conditions.
@@ III OWO Fair Learning Mechanism
where $w\!\in\![-1/u_{\min}^{\alpha},-1/u_{\max}^{\alpha}]^{K}$ are the dual conjugate variables (one per user), $(-F_{\alpha})^{\star}$ the convex conjugate of the negated fairness function, and the utilities are bounded as $u_{ti}\in[u_{\min},u_{\max}]^{K}$. Using [19, Lemma 1], this function can be written analytically as:
@@
where $g_{ti}^{\theta}\!\doteq\!\nabla_{\theta}\Psi_{ti}(w_{ti},\theta_{ti})\!=\!\!\sum_{k}w_{tik}\nabla\ell_{tik}(\theta_{ti})$. The regret of this update depends on the initial model [18, Th. 2.13]:
@@
Finally, the dual updates balance the priorities of users when training the model, by weighting the respective gradients, so as to be fair across all $m$ slots of each round. The learning of these weights is performed with strongly convex [^3] A function is $\mu$-strongly convex if $\nabla^{2}f(x)\!\succeq\!\mu I$ $\forall x$. Indeed, here holds $\nabla_{w}^{2}\Psi_{ti}(\theta,w)=diag((-w_{k}^{-\frac{\alpha+1}{\alpha}}/\alpha)_{k\in\mathcal{K}})\succeq(u_{min}^{1+1/\alpha}/\alpha)I$. OGD:
@@
Complexity and Performance Analysis. In terms of computations, the algorithm requires one outer-loop OGD update, and two first-order updates per slot within the round. Depending on the geometry of $\Theta$, these updates could be computed even in closed form. In terms of memory, unlike some prior MTL works (e.g., [16], [5]), our method does not require storing gradients separately for each user, as it directly computes the weighted gradient. For example, if the model concerns a Deep Learning network, the required gradients can be obtained through a single backpropagation on the weighted aggregated loss across all users. The performance of our scheme is stated in the next theorem (proved in appendix):
@@
Compared to a natural baseline of single-round learning (SRL), where at the begging of each round the shared model is trained from scratch, our regret bound scales as $\mathcal{O}(D_{\Theta}^{\star}/\sqrt{m})$, while SRL yields $\mathcal{O}(D_{\Theta}/\sqrt{m})$ as $T\to\infty$. Hence, Algorithm 1 achieves a multiplicative improvement by a factor $\alpha=D_{\Theta}/D_{\Theta}^{\star}$. Whenever $D_{\Theta}^{\star}\ll D_{\Theta}$, this improvement is substantial, whereas when $D_{\Theta}\approx D_{\Theta}^{\star}$, the two bounds are of the same order, and our method is only worse by a constant factor. This shows that the proposed approach is advantageous when many rounds are available and tasks are highly similar, as good initializations can then be learned.
@@ IV Performance Evaluation
Figure 3: Accumulated fairness (left) and utilities (right) across rounds. OWO-FMTL (LAST) achieves the best balance between fairness and user utilities.
Since computing $\theta_{t}^{\star}$ in this non-convex setting is impractical, we approximate it using either the parameters from the last slot (LAST) of the round or the average parameters (AVG) across the round. In terms of fairness, Fig. 3 shows that OWO-FMTL (LAST) achieves the best trade-off between fairness and user utilities across training rounds, compered to different constant weighting schemes (CWS), including the standard MTL approach that simply averages the task losses [^4] We consider constant weighting schemes as reasonable fairness benchmarks, since prior works (e.g., [8, 4]) when extended to a multi-user setting, they simply average users’ losses.. As CWS approaches rely on fixed task weights, their performance depends on whether the chosen weights align with the optimal ones. This limitation becomes evident in practice, as CWS exhibit larger fairness and utility gaps, particularly for the adversity-affected second user. In contrast, our dynamic task-weighting approach resolves this issue, achieving $\approx 20$$40\%$ higher fairness and $\approx 10$$30\%$ higher utilities across users.
Figure 4: Test losses on the validation data of each round. Test loss of SRL is around the same level as random guess, while OWO-FMTL (LAST) expedites the learning by carrying over the knowledge from previous rounds.
Regarding the impact of the outer-loop, Fig. 4 illustrates the evolution of test losses on unseen validation data across rounds. When the model is trained from scratch at each round (SRL), the limited learning horizon prevents noticeable progress. In contrast, our method exhibits a downward trend in test loss, confirming experimentally that the outer‐loop learns meaningful initializations over time, allowing the model to quickly adapt to each round’s tasks. It is worth noting that a similar result is expected in the absence of adversarial perturbations; however, the fairness–utility tension is more pronounced in highly-volatile and adversarial conditions, making the learning process even more challenging.
@@ V Conclusions and Future Work
Future mobile networks aim to deliver AI services through edge computing at the RAN, where MTL offers an efficient way to address limited resources. To this end, we introduce OWO-FMTL, a lifelong extension of MTL that effectively balances fairness and learning efficiency under both stochastic and adversarial environments. It achieves an $\mathcal{O}(1/\sqrt{m})$ vanishing round-average fair regret in the convex setting, while numerical results further demonstrate its superiority over representative baselines across diverse conditions. As future work, we plan to extend the evaluation to real-world wireless datasets, develop alternative meaningful fairness notions for this setting, and to explore different outer-loop strategies, all supported by theoretical fairness guarantees.
@@ VI Appendix
where the last residual term, under the standard perturbation-restrictiveness assumption on the environment of [19, Th. 1], tends to $\mathit{o}_{m}(1)$. Otherwise achieving vanishing horizon-fairness regret is not possible. Thus, the overall regret is:

### v2

@@ I Introduction
Still, MTL on its own has already shown significant promise across various applications. For instance, it was found to improve RAN-related tasks such as carrier prediction, resource management, modulation and signal classification, as well as radio environment and mobility characterization [17, 1, 12, 30, 9]. Further, MTL has been used for multi-objective network optimization, collaborative control and multi-task edge inference, demonstrating that shared model representations can enhance learning [22, 26]. A natural next step is to use MTL for dynamic user tasks. Online learning (OL) is well suited to this setting and fairness-aware Online Convex Optimization (OCO) has been already explored for O-RAN workload assignment and energy minimization [3, 4]. However, its integration with MTL for dynamic user-task learning remains unstudied.
@@
The operation of this system is modeled as an OWO fairness problem that we tackle with a two-layer learning algorithm based on the theory of OCO. At the higher layer (outer-loop), the system learns to initialize the model at the start of each round; where a good initialization helps the model to adapt fast to the forthcoming jobs. At the lower layer (inner-loop) the system updates the model with the feedback [^1] Feedback may come from a small test set [6, 10], direct measurements [9], or self/unsupervised metrics. the users provide after each slot’s jobs are executed. The algorithm offers optimality guarantees or, formally, zero fairness regret: it achieves the same performance as a model that would have been selected by a clairvoyant oracle, and this holds for any possible sequence of jobs and tasks. At the same time, the algorithm has minimal computing and storage requirements, which is in stark contrast to other MTL methods that calculate and store as many gradients as the tasks. Our extensive performance evaluation using both (convex) kernel regression and (non-convex) deep learning experiments, verifies the practical gains of the algorithm over several benchmarks and state-of-the-art MTL solutions.
@@ II System Model and Problem Statement
Model. We consider an AI-RAN system serving a set of users $\mathcal{K}$, who create ML tasks in a streaming fashion. The system operates in mixed-time scale with $m$ slots per round $t\!\leq\!T$. During each round $t$, every user $k\in\mathcal{K}$ needs to perform a task, denoted $(k,t)$, that involves $m$ sequential jobs (one per slot). The system supports the users through a shared model which is trained dynamically with data from all tasks. For instance, consider a mobile VM service running at the edge [11]. Each user $k$ corresponds to a mobile device with the task of object detection through a video stream. A single task $(k,t)$ then represents one usage session of the service under a particular context, e.g., lighting and weather conditions.
@@ III OWO Fair Learning Mechanism
where $w\!\in\![-1/u_{\min}^{\alpha},-1/u_{\max}^{\alpha}]^{K}$ are the dual conjugate variables (one per user), $(-F_{\alpha})^{\star}$ the convex conjugate of the negated fairness function, and the utilities are bounded as $u_{ti}\in[u_{\min},u_{\max}]^{K}$. Using [21, Lemma 1], this function can be written analytically as:
@@
where $g_{ti}^{\theta}\!\doteq\!\nabla_{\theta}\Psi_{ti}(w_{ti},\theta_{ti})\!=\!\!\sum_{k}w_{tik}\nabla\ell_{tik}(\theta_{ti})$. The regret of this update depends on the initial model [20, Th. 2.13]:
@@
Finally, the dual updates balance the priorities of users when training the model, by weighting the respective gradients, so as to be fair across all $m$ slots of each round. The learning of these weights is performed with strongly convex OGD:
@@
Complexity and Performance Analysis. In terms of computations, the algorithm requires one outer-loop OGD update, and two first-order updates per slot within the round. Depending on the geometry of $\Theta$, these updates could be computed even in closed form. In terms of scalability, unlike some prior MTL works (e.g., [18], [7]), our method avoids per-user gradient storage by directly computing the weighted gradient. For instance, if the model concerns a Deep Learning network, this is equivalent to performing a single backpropagation per iteration on the weighted aggregated loss, plus lightweight per-user scalar weighting. The performance of our scheme is stated in the next theorem (proved in appendix):
@@
Compared to a natural baseline of single-round learning (SRL), where at the beginning of each round the shared model is trained from scratch, our regret bound scales as $\mathcal{O}(D_{\Theta}^{\star}/\sqrt{m})$, while SRL yields $\mathcal{O}(D_{\Theta}/\sqrt{m})$ as $T\to\infty$. Hence, Algorithm 1 achieves a multiplicative improvement by a factor $\alpha=D_{\Theta}/D_{\Theta}^{\star}$. Whenever $D_{\Theta}^{\star}\ll D_{\Theta}$, this improvement is substantial, whereas when $D_{\Theta}\approx D_{\Theta}^{\star}$, the two bounds are of the same order, and our method is only worse by a constant factor. Thus, our theory formalizes the underlying tradeoff: the proposed approach is advantageous when many rounds are available and cross-round similarity is present, while remaining robust when tasks are diverse.
@@ IV Performance Evaluation
Figure 3: Accumulated fairness (left) and utilities (right) across rounds.
Since computing $\theta_{t}^{\star}$ in this non-convex setting is impractical, we approximate it using either the parameters from the last slot (LAST) of the round or the average parameters (AVG) across the round. In terms of fairness, Fig. 3 shows that OWO-FMTL (LAST) achieves the best trade-off between fairness and user utilities across training rounds, compared to different constant weighting schemes (CWS), including the standard MTL approach that simply averages the task losses. As CWS approaches rely on fixed task weights, their performance depends on whether the chosen weights align with the optimal ones. This limitation becomes evident in practice, as CWS exhibit larger fairness and utility gaps, particularly for the adversity-affected second user. In contrast, our dynamic task-weighting approach resolves this issue, achieving $\approx 20$$40\%$ higher fairness and $\approx 10$$30\%$ higher utilities across users.
Figure 4: Test losses on the validation data of each round.
Regarding the impact of the outer-loop, Fig. 4 illustrates the evolution of test losses on unseen validation data across rounds. When the model is trained from scratch at each round (SRL), the short horizon limits learning, keeping test loss near random guessing. In contrast, our method steadily reduces test loss, indicating that the outer-loop learns meaningful initializations that enable rapid adaptation to each round’s tasks. A similar result is expected without adversarial perturbations. However, the fairness–utility tension is more pronounced under highly volatile and adversarial conditions, making the learning process even more challenging.
@@ V Conclusions and Future Work
Future mobile networks aim to deliver AI services through edge computing at the RAN, where MTL offers an efficient way to address limited resources. To this end, we proposed OWO-FMTL, a lifelong extension of MTL that balances fairness and learning efficiency, and achieves an $\mathcal{O}(1/\sqrt{m})$ vanishing round-average fair regret in the convex setting. Numerical results further demonstrate consistent gains over representative baselines across diverse operating conditions. As future work, we plan to extend our evaluation to real-world wireless datasets to better reflect practical communication environments and constraints.
Acknowledgments
This work has been supported by the National Growth Fund through the Dutch 6G flagship project “Future Network Services”, and the European Commission through Grant No. 101139270 (ORIGAMI), 101168816 (FINALITY) and 101192462 (FLECON-6G).
@@ VI Appendix
where the last residual term, under the standard perturbation-restrictiveness assumption on the environment of [21, Th. 1], tends to $\mathit{o}_{m}(1)$. Otherwise achieving vanishing horizon-fairness regret is not possible. Thus, the overall regret is: